OpenStudy (anonymous):

MEDALLL A system of equations is shown below: 5x - 5y = 10 3x - 2y = 2 Part A: Create an equivalent system of equations by replacing one equation with the sum of that equation and a multiple of the other. Show the steps to do this. Part B: Show that the equivalent system has the same solution as the original system of equations.

OpenStudy (cwrw238):

multiply the second equation by 2 and add it to the first then solve both systems

OpenStudy (jdoe0001):

Create an equivalent system of equations by replacing one equation with the sum of that equation and a multiple of the other. <--- hint

OpenStudy (anonymous):

@jdoe0001, that is what confuses me !

OpenStudy (jdoe0001):

ok so.. pick a multiple... any value, so we could use for one of them

OpenStudy (jdoe0001):

well... I shoudln't say a "multiple" pick an integer that is

OpenStudy (anonymous):

2 ?

OpenStudy (jdoe0001):

you'd, multiply, as cwrw238 pointed out, he's using a 2 you'd multiply one of them by some value then SUM that one and the other to get a resultant equation

OpenStudy (jdoe0001):

it could be any value, doesn't have to be 2 so... pick one

OpenStudy (anonymous):

i'll do 2, but, 3x - 2y = 2 how would you do that ?

OpenStudy (anonymous):

?

OpenStudy (jdoe0001):

ok let us use 2 then and let's multiply say the 2nd one so \(\large { \begin{array}{llll} 5x - 5y = 10&\implies &5x - 5y = 10\\ 3x - 2y = 2&\implies {\color{brown}{ \times 2}}&6x-4y=4 \\\hline\\ &&\square ?\quad \square ?= \square \end{array} }\)

OpenStudy (anonymous):

you would subtract the last things ?

OpenStudy (jdoe0001):

well.. you'd add/subtract them both vertically like you'd any values

OpenStudy (anonymous):

so, it would be 11x - 9y = 14 ?

OpenStudy (jdoe0001):

hmm lemme fix a typo there

OpenStudy (jdoe0001):

thus \(\large { \begin{array}{llll} 5x - 5y = 10&\implies &5x - 5y = 10\\ 3x - 2y = 2&\implies {\color{brown}{ \times 2}}&6x-4y=4 \\\hline\\ &&{\color{blue}{ 11x -9y=14}} \end{array} \\ \quad \\ \begin{array}{llll} 5x - 5y = 10\\ 3x - 2y = 2 \end{array}\quad \equiv\quad \begin{array}{llll} {\color{blue}{ 11x -9y=14}}\\ 3x - 2y = 2 \end{array}\quad \equiv\quad \begin{array}{llll} 5x - 5y = 10\\ {\color{blue}{ 11x -9y=14}} \end{array} }\)

OpenStudy (jdoe0001):

part B means you have to find "x" and "y" in the original and then find "x" and "y" in the new system of equations to show that both match

OpenStudy (anonymous):

so how do you do that? plug in 0 ?

OpenStudy (jdoe0001):

well... use substitutions or elimination you'd sure have covered that

OpenStudy (anonymous):

okay, thank you :)

OpenStudy (jdoe0001):

yw